There are 132 more silver dollars than 50-cent pieces.
The ratio of the silver dollars to 50-cent pieces is 5:2.
So how many of each coin are there?

Answer by stanbon(75887)

There are 132 more silver dollars than 50-cent pieces.
Let # of 50-cent pieces by "x"; # of silver dollars is "x+132"
------------------------
The ratio of the silver dollars to 50-cent pieces is 5:2.
So how many of each coin are there?
EQUATION:
(x+132)/x = 5/2
Cross-multiply to get:
2x+264 = 5x
3x = 264
x = 88 (# of 50-cent pieces)
x + 132 = 220 (# of silver dollars)
--------------------
Cheers,
Stan H.